2. List the Source of Variance (SV) and their corresponding degress of freedom (df) for each of the following ANOVA.
a. n=3, J=5

b. n=5, J=20

c. n=2, J=10
Answer #1

Ans:

a)number of treatments=n=3

Number of subjects in each group=5

Total number of subjects=3*5=15

df(between)=3-1=2

df(within)=15-3=12

2)number of treatments=n=5

Number of subjects in each group=20

Total number of subjects=5*20=100

df(between)=5-1=4

df(within)=100-5=95

c)number of treatments=n=2

Number of subjects in each group=10

Total number of subjects=2*10=20

df(between)=2-1=1

df(within)=20-2=18